OpenStudy (anonymous):

Write the following expression as the sum of three partial fractions:

OpenStudy (anonymous):

\[E = \frac{ 2x - 5x^2 }{ (x+1)^3 }\]

sam (.sam.):

Hmm try \[\frac{A}{x+1}+\frac{B}{(x+1)^2}+\frac{C}{(x+1)^3}\]

OpenStudy (anonymous):

Ok lemme try that and get back to you.

sam (.sam.):

Alright

OpenStudy (anonymous):

but won't I have to take x as -1 all the time?

OpenStudy (anonymous):

But then A and B and C will always come as 0. Can I take any value for x and then solve them simultaneously??

sam (.sam.):

You multiply the whole thing by \((x+1)^3\), you get \[2x-5x^2=A(x+1)^2+B(x+1)+C\]

sam (.sam.):

\[ \frac{ 2x - 5x^2 }{ (x+1)^3 }= \frac{A}{x+1}+\frac{B}{(x+1)^2}+\frac{C}{(x+1)^3} \\ \\ 2x-5x^2=A(x+1)^2+B(x+1)+C\]

OpenStudy (anonymous):

Wait wait hold on. Does the equation become A(x+1)^2 + B(x+!) + C = 2x - 5x^2

sam (.sam.):

Yes correct

sam (.sam.):

Then maybe we'll comapare coefficients for this, \[x^2:~~~ -5=A \\ \\ x:~~~2=2A+B \\ \\ Constant:~~~0=A+B+C\]

OpenStudy (anonymous):

Wait, we can do that?? I thought we need to put in values for x and then solve.

sam (.sam.):

You can compare them straight away

OpenStudy (anonymous):

So far i've got C = -1 and A+B = 7

OpenStudy (anonymous):

But you can't compare them, on the left it's (x+1)^2 and on the right it's just x^2. That won't make sense, would it.

OpenStudy (anonymous):

Oh but wait, you accounted for that. Sorry :P

sam (.sam.):

How dow you get that? \[-5x^2=Ax^2 \\ \\ 2x=2Ax+Bx \\ \\ 0=A+B+C\]

sam (.sam.):

Yeah

sam (.sam.):

A will be -5 Then you can find B and C already

OpenStudy (anonymous):

Yup, i've got 12 and -7.

sam (.sam.):

correct

hartnn (hartnn):

just a note : you can also *mix* the 2 methods, if you put x=-1, you directly get C then go for comparison :)

OpenStudy (anonymous):

Ok, and the next pert of the question says check that you've got it right. How do I do that??

sam (.sam.):

Aw

OpenStudy (anonymous):

Aw?? :P

OpenStudy (anonymous):

*part.

sam (.sam.):

I'll make common denominators for that :P, but I'm sure there's another method, just can't think of it right now

OpenStudy (anonymous):

Meh. @hartnn , got anything?

hartnn (hartnn):

common denominator is the best(and maybe the only) way

OpenStudy (anonymous):

So i make the partial fractions into one fraction again??

sam (.sam.):

Yes

OpenStudy (anonymous):

Dear lord. Why does my math teacher do this to me, WHYYY

OpenStudy (anonymous):

like, he makes me expand in into three, and then says make it one again ughh

OpenStudy (agent0smith):

That's how you check your work to make sure you expanded it correctly :P It is tedious, though.

OpenStudy (anonymous):

Ikr??